The tune-up specifications of a car call for the spark plugs to be tightened to a torque of 30 N⋅m . You plan to tighten the plugs by pulling on the end of a 21-cm-long wrench. Because of the cramped space under the hood, you'll need to pull at an angle of 125 ∘ with respect to the wrench shaft.
Answer #1

The OP ahs forgotten to write Solve for Force :

here

T = 30 N.m
S = 21 cm = .21m
Net angle wrench = 125 - 90 = 35 degrees

As we know ,

Torque = Force * perpendicular Distance
T = F * 0.21*Cos35
F = 30 / (0.21*Cos35)
F = 174.39 N